# The origin of supermassive black holes at cosmic dawn

### v2

@@ Abstract
quasars: supermassive black hole – galaxies: formation – galaxies: haloes – accretion
pubyear: 2015pagerange: The origin of supermassive black holes at cosmic dawnA
@@ 1 Introduction
Black holes are ubiquitous in the Universe. A plethora of stellar black holes dot galaxies, and they formed as an end state of massive stars. However, another class of black holes occupy the centre of galaxies with masses at least a million times that of the Sun and are known as the supermassive black holes (SMBHs) (e.g. Kormendy & Richstone, 1995; Richstone et al., 1998). The SMBHs at low redshifts can be explained by the accretion of matter onto a seed black hole over the lifetime of a galaxy (e.g. Soltan, 1982; Yu & Tremaine, 2002; Shankar et al., 2004). However, the existence of supermassive black holes in the early Universe at redshifts similar to when the first galaxies originated (Mortlock et al., 2011; Larson et al., 2023) poses a challenge to theoretical models (e.g. Volonteri et al., 2021).
@@
Although the super-Eddington accretion is observed in some X-ray binaries (e.g. Okuda, 2002) and some narrow-line Seyfert-I galaxies (Mineshige et al., 2000; Collin & Kawaguchi, 2004; Du et al., 2014), but the consensus is that the super-Eddington accretion is prone to radiation feedback and outflows, hence it may not be sustained. Therefore, alternative mechanisms have been proposed to explain the high-z SMBHs. These range from SMBHs beginning growth from primordial black holes (e.g. Kawasaki et al., 2012) to mechanisms invoking direct collapse black holes (e.g. Bromm & Loeb, 2003). Some studies indicate that super-Eddington accretion may operate in brief episodes (Johnson & Bromm, 2007; Park & Ricotti, 2011; Tang et al., 2019) or that the outflows might not be triggered until the black hole reaches a critical mass (e.g. King, 2003; Li, 2012). The recent detection of SMBHs at very high redshifts (e.g. at $z\approx 8.7$ by Larson et al. 2023) has brought this debate to the forefront (see also Maiolino et al. 2023a, b; Kokorev et al. 2023). The exact mechanism of the accretion in the vicinity of a black hole and the process through which the radiation emerges, and couples with the accretion flow needs to be understood.
@@
In SMBHs as well, the disc likely regulates the efficiency of the radiation feedback and, hence, the accretion rate into the hole. The efficiency is hotly debated. According to some studies, the value is around 0.1, or lower due to the photon trapping in the inner disc (Begelman, 1979). Hence, considering that roughly 10% of the accretion energy is radiated (Rees, 1984; Yu & Tremaine, 2002; Ueda et al., 2003), and if the accretion rate is high enough, this radiation may stop the accretion and drive outflows. However, for outflows to occur, a minimum black hole mass is required (Silk & Rees, 1998; King, 2003).
@@
The standard Bondi’s mass accretion rate used in cosmological simulation is for accretion without angular momentum. However, the accretion likely has angular momentum, forming a disc near the black hole that regulates the mass inflow rate and the growth of the black hole (e.g. Rosas-Guevara et al., 2015). Another uncertainty is about the fraction of mass inflow entering the halo that finds its way to the hole and the fraction that results in the formation of the galaxy and its components (e.g. McAlpine et al., 2017; Bower et al., 2017). A primary question in this regard is how the Bondi accretion solution and the mass accretion rate formula are modified when considering a steady spherical flow into the combined potential of the hole and the halo. The subsequent task is to put this in the context of the gas viriallization and cooling in a halo and the formation of galaxy in-situ with SMBH (Dubois et al., 2015; Rosas-Guevara et al., 2016; Anglés-Alcázar et al., 2017; Habouzit et al., 2017; Bower et al., 2017; McAlpine et al., 2017, 2018; Weinberger et al., 2018).
@@
In this paper, we address the problem of accretion onto a black hole at the centre of a dark matter halo. We consider both a Hernquist and an NFW halo profile. We focus on bringing out the physics of an isothermal steady flow in the combined potential of the black hole and the halo. We derive the critical points of the flow and investigate the possibility of trans-sonic accretion, and the role of the black hole.
We also investigate the consequences of our solutions for the growth of SMBHs in early haloes and then derive the evolution of the Eddington ratio, to investigate the need for super-Eddington accretion to explain the high redshift SMBHs, detected by JWST and earlier studies (Fan et al., 2001, 2003; Mortlock et al., 2011; Inayoshi et al., 2020; Larson et al., 2023).
@@ 2 Spherically symmetric accretion into dark matter haloes
The equations for the mass and the momentum conservation for a spherically-symmetric steady flow are,
@@
The sound speed, $c_{s}=\sqrt{{\rm k_{B}}T/(\mu{\rm m_{H}}})$, where $\mu$ is the mean molecular weight, $m_{\rm H}$ is the mass of the hydrogen atom, and $T$ is the temperature which is constant for an isothermal flow, therefore Eq 3 can be readily solved to obtain,
@@ 2.1 Accretion into the centre of a dark matter halo
In this section we consider an isothermal accretion flow into a dark matter halo. We first consider the Hernquist profile for the halo (Hernquist, 1990). The potential for a Hernquist dark matter halo is,
@@
Figure 1: The family of solutions for the steady spherical flow to the centre of a Hernquist dark matter halo, obtained by varying $C$ in Eq. 7; showing Mach number, $\mathcal{M}$, as a function of the dimensionless radial distance, $x=r/l_{\rm h}$, where $l_{h}=GM_{\rm h}/c_{\rm s}^{2}$. Top panel: when no black hole is present at the centre ($\beta=M_{\rm bh}/M_{\rm h}=0$), there are two critical points, and the trans-sonic flow does not go all the way to the centre. When a black hole is present then there are three critical points, and the trans-sonic flow into the centre is possible either through the innermost critical point (bottom panel) or through the outermost critical point (middle panel) depending on the value of $\alpha=a/l_{\rm h}$. The bottom panel is for $\beta=10^{-6}$ and $\alpha=10^{-1}$. The middle panel is for the same $\beta=10^{-6}$ but for a lower $\alpha=10^{-2}$. The location of the critical points depend on the values of $\alpha$ and $\beta$, and can also be visualised in Fig. 2.
The family of solutions is shown in Fig. 1 top panel, by varying the value of $C$. The flow can be analysed by the knowledge of critical points. The location of the critical points of the flow ($r_{c}$) can be determined by setting the right hand side of Eq. 6 to zero, to obtain,
@@
where, $x_{c}=r_{c}/l_{h}$, and, $\alpha=a/l_{\rm h}$. Two critical points are possible for $\alpha<1/8$, one is located at $r_{c}<a$, the other at $r_{c}>a$ (see red curve in Fig. 2). The two critical points approach each other as $a$ increases or $l_{\rm h}$ decreases, and for $\alpha=a/l_{\rm h}=1/8$, they merge into a single critical point located at $r_{c}=a$. No critical point exists for $\alpha>1/8$.
@@
where we have used $\mu=0.6$. We note that that this maximum temperature is roughly same as the virial temperature of the halo. In fact, an analogous calculation for an NFW halo with concentration parameter $c\approx 5$ would yield $T_{\rm max}\sim T_{\rm vir}$ with $\alpha$ in that case is defined as $\alpha=R_{\rm s}/l_{\rm h}$ where $R_{s}$ is the NFW scale radius related to the virial radius as $R_{s}=R_{\rm vir}/c$.
For $\alpha$ less than a critical value, which is $1/8$ for the Hernquist halo (and approximately $0.1$ for the NFW halo with $c=5$), the flow can proceed along the solution through the outer critical point, however, such a solution is unphysical as it circles around the inner critical point and is double valued (Fig. 1 top panel). Therefore, trans-sonic accretion all the way to the centre is not possible for a Hernquist or an NFW dark matter halo. However, this outcome changes when a seed black hole is present as we investigate in the next section.
@@ 2.2 Accretion onto a seed black hole in a Hernquist dark matter halo
In this section we consider the effect of the presence of a seed black hole at the centre of a Hernquist halo. The combined potential of a Hernquist dark matter halo and the black hole at its centre is,
@@
where we have defined the ratio of the black hole mass to the halo as, $\beta=M_{\rm bh}/M_{\rm h}$. The flow solution for this total potential is,
@@
In Fig. 2 we have shown the locus of critical points as a function of $\alpha$ for four different values of $\beta$. The location of the critical point for a specific $\alpha$ and $\beta$ can be read from the appropriate curve.
In general there are three critical points for a specific $\alpha$ when a black hole is present (black, magenta and orange curves) compared to only two when there is no black hole (red curve). A trans-sonic flow into the black hole can proceed through the innermost or the outermost critical point.
The trans-sonic flow through the innermost critical point is shown in Fig. 1 bottom panel, and this solution is akin to the standard Bondi accretion onto a black hole. The trans-sonic flow through the outermost critical point (in Fig. 1 middle panel) is due to the halo. There are unphysical solutions as well shown in Fig. 1. For example, in the middle panel, the solution shown by the red curve passes through the innermost critical point, but then it circles around the intermediate critical point and turns back, hence it is double-valued and unphysical. The corresponding values of $\alpha$ and $\beta$ for which such unphysical solutions occur are shown as dashed part of the curves in Fig. 2).
Figure 2: Top panel: Location of the critical points for accretion into the combined potential of a black hole and a Hernquist halo, shown as a function of $\alpha=a/l_{\rm h}$, for $\beta=0$ (red), $10^{-9}$ (black), $10^{-7}$ (magenta) and $10^{-5}$ (yellow). The critical points are the points at which the vertical lines corresponding to a constant $\alpha$ intersects the curves corresponding to a particular $\beta$. For $\beta=0$ there are a maximum of two critical points, and for $\beta\neq 0$ a maximum of three critical points. Further, for non-zero $\beta$, the trans-sonic flow to the centre may proceed through the innermost or the outermost critical point. For high $\alpha$, the solution proceeds through the innermost critical point, however, as $\alpha$ decreases below a transition value (shown as filled circles) which depends on $\beta$ (Eq. 19 for Hernquist and Eq. 24 for NFW), the solution through the innermost critical point is unphysical (shown by the dashed portion of the curves and also see the red curve in the middle panel of Fig. 1)
, then the other trans-sonic solution through the outermost critical point takes-over. Bottom panel: same as the top panel but for an NFW halo potential with concentration parameter $c=5$.
The two types of solutions operate in two separate regimes of $\alpha$. Also, the mass accretion rate for the two types of solutions are significantly different. The mass accretion rate can be calculated from Eq. 1 and 2; then using $v=c_{s}$ at the critical point $r=r_{c}$; and $\rho=\rho_{\infty}$, $v=0$ at infinity. The mass accretion rate is thus given by,
@@
For the combined potential of a Hernquist dark matter halo and a black hole (Eq. 10) we obtain,
@@
where $\lambda$ is the dimensionless mass accretion rate. If the location of the critical point is known, we can obtain the exact value of the mass accretion rate. In two special cases the critical point can be recovered analytically.
The innermost critical point can be recovered by setting the limit $x_{c}\ll\alpha$ in Eq. 13, that yields $x_{c}\approx\beta/2$. For this the dimensionless mass accretion rate (from Eq. 15) is given by,
@@
For high values of $\alpha$, $\lambda\approx 1.12\beta^{2}$, that is the same as for the standard Bondi accretion. This is not surprising since the innermost critical point is due to the black hole, and flow through it approaches the standard Bondi accretion.
@@
The critical point is then given by, $x_{\rm c}\approx\big{(}1+\beta\big{)}/2$. Substituting the critical point in Eq. 15 together with, $\alpha/x_{c}\approx 0$, yields the following expression for the dimensionless mass accretion rate,
@@
As $\beta\ll 1$, we have $\lambda\approx 1.12$. This also appears similar to the standard Bondi accretion however the difference is that the mass accretion rate is proportional to $M_{\rm h}^{2}$ instead of $M_{\rm bh}^{2}$ (see Eq. 15).
The two trans-sonic solutions are separated by a transition $\alpha$, and for the transition $\alpha$ the flow solution passes through both the innermost and the outermost critical point.
The transition $\alpha$ depends on the value of $\beta$, and can be derived by imposing the condition that the trans-sonic flow passes through both the outermost and the innermost critical point (see Appendix A). The transition $\alpha$ thus recovered is,
@@
For $\alpha\gg\alpha_{\rm T}$, the mass accretion rate is low and it approaches the standard Bondi’s mass accretion rate (Bondi, 1952). For $\alpha\ll\alpha_{\rm T}$, the mass accretion rate becomes high ($\lambda\approx 1.12$ or $\dot{M}_{\rm bh}\propto M_{\rm h}^{2}$) as it is controlled by the halo (Fig. 4).
The value of $\alpha$ decreases with redshift as the halo mass increases (Correa et al., 2015a), that indicates two phases of accretion, beginning with a phase of slow growth initially for a high $\alpha\gg\alpha_{\rm T}$, and then a phase of rapid growth at $\alpha$ below $\alpha_{\rm T}$ (Eq. 19).
3 Accretion onto a seed black hole in a NFW dark matter halo
Figure 3: The mach number, $\mathcal{M}$, as a function of $x=r/l_{\rm h}$ for accretion onto the black hole residing in dark matter halo shown by the curves that rise towards the centre; for a Hernquist halo (blue curves) and a NFW halo (red curves). For NFW haloes, we have chosen a concentration parameter, $c=5$, that is appropriate for high redshift haloes (e.g. Prada et al., 2012; Correa et al., 2015b). In general, there are three critical points in each case. Top panel: for the trans-sonic flow through the outer-most critical point (that occurs for $\alpha=0.01,\ \beta=10^{-6}$). Bottom panel: for the flow through the inner-most critical point (for $\alpha=0.1$ and $\beta=10^{-6}$). For the Hernquist halo $\alpha=a/l_{h}$, and for the NFW halo, $\alpha=R_{\rm s}/l_{\rm h}$, where $R_{\rm s}$ is the scale-radius that is related to the virial (truncation) radius as $R_{\rm s}=R_{\rm vir}/c$. The paramter $l_{\rm h}=GM_{\rm h}/c_{s}^{2}$.
Figure 4: The dimensionless mass accretion rate, $\lambda$, as a function of $\alpha$, for $\beta=10^{-9}$ (red), $\beta=10^{-7}$ (black) and $\beta=10^{-5}$ (blue); for a Hernquist halo (dotted curves), and for an NFW halo (dashed curve). Also, the parameter $\alpha=a/l_{\rm h}$ for the Hernquist halo, and $\alpha=R_{s}/l_{\rm h}$ for the NFW halo, where $R_{s}$ is the NFW scale radius, related to the virial radius as $R_{s}=R_{\rm vir}/c$, and we take $c=5$. The dashed portion of the curves shows the mass accretion rate when the trans-sonic solution is through the innermost critical point (that occurs for a higher $\alpha$). The solid portion of the curve shows the mass accretion rate when the flow proceeds through the outermost critical point (for smaller $\alpha$). $\lambda$ increases from a low value ($\approx\beta^{2}$) with decreasing $\alpha$ and becomes constant ($\lambda\approx 1.12$), when the trans-sonic solution through the outermost critical point takes-over.

### v3

@@ Abstract
quasars: supermassive black hole – galaxies: formation – galaxies: haloes – accretion, accretion discs
pubyear: 2024pagerange: The origin of supermassive black holes at cosmic dawnA
@@ 1 Introduction
Black holes are ubiquitous in the Universe. A plethora of stellar black holes dot galaxies, and they formed as an end state of massive stars. However, another class of black holes occupy the centres of galaxies with masses at least a million times that of the Sun and are known as the supermassive black holes (SMBHs) (e.g. Kormendy & Richstone, 1995; Richstone et al., 1998). The SMBHs at low redshifts can be explained by the accretion of matter onto a seed black hole over the lifetime of a galaxy (e.g. Soltan, 1982; Yu & Tremaine, 2002; Shankar et al., 2004). However, the existence of supermassive black holes in the early Universe at redshifts similar to when the first galaxies originated (Mortlock et al., 2011; Larson et al., 2023) poses a challenge to theoretical models (e.g. Volonteri et al., 2021).
@@
Although the super-Eddington accretion is observed in some X-ray binaries (e.g. Okuda, 2002) and some narrow-line Seyfert-I galaxies (Mineshige et al., 2000; Collin & Kawaguchi, 2004; Du et al., 2014); however the consensus is that the super-Eddington accretion is prone to radiation feedback and outflows, hence it may not be sustained. Therefore, alternative mechanisms have been proposed to explain the high-z SMBHs. These range from SMBHs beginning growth from primordial black holes (e.g. Kawasaki et al., 2012) to mechanisms invoking direct collapse black holes (e.g. Bromm & Loeb, 2003). Some studies indicate that super-Eddington accretion may operate in brief episodes (Johnson & Bromm, 2007; Park & Ricotti, 2011; Tang et al., 2019) or that the outflows might not be triggered until the black hole reaches a critical mass (e.g. King, 2003; Li, 2012). The recent detection of SMBHs at very high redshifts (e.g. at $z\approx 8.7$ by Larson et al. 2023) has brought this debate to the forefront (see also Maiolino et al. 2023a, b; Kokorev et al. 2023). The exact mechanism of the accretion in the vicinity of a black hole, the resulting production of radiation, and its coupling with the accretion flow need to be understood.
@@
In SMBHs as well, the disc likely regulates the efficiency of the radiation feedback and, hence, the accretion rate into the hole. The efficiency is hotly debated. According to some studies, the value is around $0.1$ or lower due to the photon trapping in the inner disc (Begelman, 1979). Hence, considering that roughly $10$% of the accretion energy is radiated (Rees, 1984; Yu & Tremaine, 2002; Ueda et al., 2003), and if the accretion rate is high enough, this radiation may stop the accretion and drive outflows. However, for outflows to occur, a minimum black hole mass is required (Silk & Rees, 1998; King, 2003).
@@
The standard Bondi’s mass accretion rate used in cosmological simulation is for accretion without angular momentum. However, the accretion likely has angular momentum, forming a disc near the black hole that regulates the mass inflow rate and the growth of the black hole (e.g. Rosas-Guevara et al., 2015). Another uncertainty is about the fraction of mass inflow entering the halo that finds its way to the hole and the fraction that results in the formation of the galaxy and its components (e.g. McAlpine et al., 2017; Bower et al., 2017). A primary question in this regard is how the Bondi accretion solution and the mass accretion rate formula are modified when considering a steady spherical flow into the combined potential of the hole and the halo. The subsequent task is to put this in the context of the gas virialization and cooling in a halo and the formation of galaxy in-situ with SMBH (Dubois et al., 2015; Rosas-Guevara et al., 2016; Anglés-Alcázar et al., 2017; Habouzit et al., 2017; Bower et al., 2017; McAlpine et al., 2017, 2018; Weinberger et al., 2018).
@@
In this paper, we address the problem of accretion onto a black hole at the centre of a dark matter halo. We consider both a Hernquist and an NFW halo profile. We focus on bringing out the physics of an isothermal steady flow in the combined potential of the black hole and the halo. We derive the critical points of the flow, determine the possibility of trans-sonic accretion, and investigate the role of the black hole.
We also study the consequences of our solutions for the growth of SMBHs in early haloes and then derive the evolution of the Eddington ratio to investigate the need for super-Eddington accretion to explain the high redshift SMBHs detected by JWST and earlier studies (Fan et al., 2001, 2003; Mortlock et al., 2011; Inayoshi et al., 2020; Larson et al., 2023).
@@ 2 Spherically symmetric accretion into dark matter haloes
The equations for the mass and the momentum conservation for a spherically symmetric steady flow are,
@@
The sound speed, $c_{s}=\sqrt{{\rm k_{B}}T/(\mu{\rm m_{H}}})$, where $\mu$ is the mean molecular weight, $m_{\rm H}$ is the mass of the hydrogen atom, and $T$ is the temperature which is constant for an isothermal flow. Therefore Eq 3 can be readily solved to obtain,
@@ 2.1 Accretion into the centre of a dark matter halo
In this section, we consider an isothermal accretion flow into a dark matter halo. We first consider the Hernquist profile for the halo (Hernquist, 1990). The potential for a Hernquist dark matter halo is,
@@
Figure 1: The family of solutions for the steady spherical flow to the centre of a Hernquist dark matter halo, obtained by varying $C$ in Eq. 7; showing Mach number, $\mathcal{M}$, as a function of the dimensionless radial distance, $x=r/l_{\rm h}$, where $l_{h}=GM_{\rm h}/c_{\rm s}^{2}$. Top panel: if no black hole is present at the centre ($\beta=M_{\rm bh}/M_{\rm h}=0$), there are two critical points, and the trans-sonic flow does not go all the way to the centre. If a black hole is present, then there are three critical points, and the trans-sonic flow into the centre is possible either through the innermost critical point (bottom panel) or through the outermost critical point (middle panel) depending on the value of $\alpha=a/l_{\rm h}$. The bottom panel is for $\beta=10^{-6}$ and $\alpha=10^{-1}$. The middle panel is for the same $\beta=10^{-6}$ but for a lower $\alpha=10^{-2}$. The location of the critical points depends on the values of $\alpha$ and $\beta$, and can also be visualised in Fig. 2.
The family of solutions is shown in Fig. 1 top panel by varying the value of $C$. The flow can be analysed by the knowledge of critical points. The location of the critical points of the flow ($r_{c}$) can be determined by setting the right-hand side of Eq. 6 to zero, to obtain,
@@
where, $x_{c}=r_{c}/l_{h}$, and, $\alpha=a/l_{\rm h}$. Two critical points are possible for $\alpha<1/8$, one is located at $r_{c}<a$, the other at $r_{c}>a$ (see red curve in Fig. 2). The two critical points approach each other as $a$ increases or $l_{\rm h}$ decreases, and for $\alpha=a/l_{\rm h}=1/8$, they merge into a single critical point at $r_{c}=a$. No critical point exists for $\alpha>1/8$.
@@
where we have used $\mu=0.6$. We note that this maximum temperature is roughly the same as the virial temperature of the halo. An analogous calculation for an NFW halo with concentration parameter $c\approx 5$ would yield $T_{\rm max}\sim T_{\rm vir}$ with $\alpha$ in that case is defined as $\alpha=R_{\rm s}/l_{\rm h}$ where $R_{s}$ is the NFW scale radius related to the virial radius as $R_{s}=R_{\rm vir}/c$.
For $\alpha$ less than a critical value, which is $1/8$ for the Hernquist halo (and approximately $0.1$ for the NFW halo with $c=5$), the flow can proceed along the solution through the outer critical point; however, such a solution is unphysical as it loops around the inner critical point and is double valued (Fig. 1 top panel). Therefore, trans-sonic accretion to the centre of a Hernquist or an NFW dark matter halo is impossible. However, this outcome changes when a seed black hole is present as we investigate in the next section.
@@ 2.2 Accretion onto a seed black hole in a Hernquist dark matter halo
In this section, we consider the effect of the presence of a seed black hole at the centre of a Hernquist halo. The combined potential of a Hernquist dark matter halo and the black hole at its centre is,
@@
where we have defined the ratio of the black hole mass to the halo as $\beta=M_{\rm bh}/M_{\rm h}$. The flow solution for this total potential is,
@@
In Fig. 2, we have shown the locus of critical points as a function of $\alpha$ for four different values of $\beta$. The location of the critical point for a specific $\alpha$ and $\beta$ can be read from the appropriate curve.
In general, there are three critical points for a specific $\alpha$ when a black hole is present (black, magenta, and orange curves) compared to only two when there is no black hole (red curve). A trans-sonic flow into the black hole can proceed through the innermost or the outermost critical point.
The trans-sonic flow through the innermost critical point is shown in Fig. 1 bottom panel, and this solution is akin to the standard Bondi accretion onto a black hole. The trans-sonic flow through the outermost critical point (in Fig. 1 middle panel) is due to the halo. There are unphysical solutions as well, shown in Fig. 1. For example, in the middle panel, the solution shown by the red curve passes through the innermost critical point, but then it loops around the intermediate critical point and turns back; hence, it is double-valued and unphysical. The corresponding values of $\alpha$ and $\beta$ for which such unphysical solutions occur are shown as dashed part of the curves in Fig. 2).
Figure 2: Top panel: Location of the critical points for accretion into the combined potential of a black hole and a Hernquist halo, shown as a function of $\alpha=a/l_{\rm h}$, for $\beta=0$ (red), $10^{-9}$ (black), $10^{-7}$ (magenta) and $10^{-5}$ (yellow). The critical points are the points at which the vertical lines corresponding to a constant $\alpha$ intersect the curves corresponding to a particular $\beta$. For $\beta=0$, there are a maximum of two critical points, and for $\beta\neq 0$, a maximum of three critical points. Further, for non-zero $\beta$, the trans-sonic flow to the centre may proceed through the innermost or the outermost critical point. For high $\alpha$, the solution proceeds through the innermost critical point; however, as $\alpha$ decreases below a transition value (shown as filled circles) which depends on $\beta$ (Eq. 19 for Hernquist and Eq. 24 for NFW), the solution through the innermost critical point is unphysical (shown by the dashed portion of the curves and also see the red curve in the middle panel of Fig. 1), then the other trans-sonic solution through the outermost critical point takes-over. Bottom panel: same as the top panel but for an NFW halo potential with concentration parameter $c=5$.
The two types of solutions operate in two separate regimes of $\alpha$. Also, the mass accretion rates for the two types of solutions are significantly different. The mass accretion rate can be calculated from Eq. 1 and 2; then using $v=c_{s}$ at the critical point $r=r_{c}$; and $\rho=\rho_{\infty}$, $v=0$ at infinity. The mass accretion rate is thus given by,
@@
For the combined potential of a Hernquist dark matter halo and a black hole (Eq. 10), we obtain,
@@
where $\lambda$ is the dimensionless mass accretion rate. If the location of the critical point is known, we can obtain the exact value of the mass accretion rate. The innermost and outermost critical points can be recovered analytically.
The innermost critical point can be recovered by setting the limit $x_{c}\ll\alpha$ in Eq. 13, which yields $x_{c}\approx\beta/2$. For this, the dimensionless mass accretion rate (from Eq. 15) is given by,
@@
For high values of $\alpha$, $\lambda\approx 1.12\beta^{2}$, the same as for the standard Bondi accretion, which is not surprising since the innermost critical point is due to the black hole, and the flow through it is akin to the standard Bondi accretion on to a point mass.
@@
The critical point is then given by, $x_{\rm c}\approx\big{(}1+\beta\big{)}/2$. Substituting the critical point in Eq. 15 together with $\alpha/x_{c}\approx 0$ yields the following expression for the dimensionless mass accretion rate,
@@
As $\beta\ll 1$, we have $\lambda\approx 1.12$, which also seems similar to the standard Bondi accretion; however, there is a subtle difference that the mass accretion rate is proportional to $M_{\rm h}^{2}$ instead of $M_{\rm bh}^{2}$ (see Eq. 15).
The two trans-sonic solutions are separated by a transition $\alpha$, and for the transition $\alpha$, the flow solution passes through both the innermost and the outermost critical point.
The transition $\alpha$ depends on the value of $\beta$ and can be derived by imposing the condition that the trans-sonic flow passes through both the outermost and the innermost critical point (see Appendix A). The transition $\alpha$ thus recovered is,
@@
For $\alpha\gg\alpha_{\rm T}$, the mass accretion rate is low, and it approaches the standard Bondi’s mass accretion rate (Bondi, 1952). For $\alpha\ll\alpha_{\rm T}$, the mass accretion rate becomes high ($\lambda\approx 1.12$ or $\dot{M}_{\rm bh}\propto M_{\rm h}^{2}$) as it is controlled by the halo (Fig. 4).
The value of $\alpha$ decreases with redshift as the halo mass increases (Correa et al., 2015a), indicating two phases of accretion, beginning with a phase of slow growth initially for a high $\alpha\gg\alpha_{\rm T}$, and then a phase of rapid growth at $\alpha$ below $\alpha_{\rm T}$ (Eq. 19).
3 Accretion onto a seed black hole in an NFW dark matter halo
Figure 3: The Mach number, $\mathcal{M}$, as a function of $x=r/l_{\rm h}$ for accretion onto the black hole residing in dark matter halo shown by the curves that rise towards the centre; for a Hernquist halo (blue curves) and an NFW halo (red curves). For NFW haloes, we have chosen a concentration parameter, $c=5$, that is appropriate for high redshift haloes (e.g. Prada et al., 2012; Correa et al., 2015b). In general, there are three critical points in each case. Top panel: for the trans-sonic flow through the outermost critical point (that occurs for $\alpha=0.01,\ \beta=10^{-6}$). Bottom panel: for the flow through the innermost critical point (for $\alpha=0.1$ and $\beta=10^{-6}$). For the Hernquist halo $\alpha=a/l_{h}$, and for the NFW halo, $\alpha=R_{\rm s}/l_{\rm h}$, where $R_{\rm s}$ is the scale-radius that is related to the virial (truncation) radius as $R_{\rm s}=R_{\rm vir}/c$. The paramter $l_{\rm h}=GM_{\rm h}/c_{s}^{2}$.
Figure 4: The dimensionless mass accretion rate, $\lambda$, as a function of $\alpha$, for $\beta=10^{-9}$ (red), $\beta=10^{-7}$ (black) and $\beta=10^{-5}$ (blue); for a Hernquist halo (dotted curves), and an NFW halo (dashed curve). Also, the parameter $\alpha=a/l_{\rm h}$ for the Hernquist halo, and $\alpha=R_{s}/l_{\rm h}$ for the NFW halo, where $R_{s}$ is the NFW scale radius, related to the virial radius as $R_{s}=R_{\rm vir}/c$, and we take $c=5$. The dashed portion of the curves shows the mass accretion rate when the trans-sonic solution is through the innermost critical point (that occurs for a higher $\alpha$). The solid portion of the curve shows the mass accretion rate when the flow proceeds through the outermost critical point (for smaller $\alpha$). $\lambda$ increases from a low value ($\approx\beta^{2}$) with decreasing $\alpha$ and becomes constant ($\lambda\approx 1.12$) when the trans-sonic solution through the outermost critical point takes over.